OpenStudy (anonymous):

Please Help I’ll Give Medal and Fan U!!!!!!! The area of a rectangle piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show steps of our work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (blake57roger):

Here are two equations from the information given: L-5 = x+1 [because it's a square] Area= 280= L * X. [because Area is length times width] Let x be the width Let L be the length Take the first equation, set it equal to L, then substitute it into the second equation, now solve for X

OpenStudy (blake57roger):

bingo theres how you get your awnser

OpenStudy (anonymous):

So which is part a and part b

OpenStudy (anonymous):

what do u mean by set it equal to L

OpenStudy (blake57roger):

one sec

OpenStudy (blake57roger):

@iPwnBunnies

OpenStudy (blake57roger):